how do i calculate 2 4/5-5 1/4
Answer by Alan3354(69443)

2 4/5-5 1/4
= 14/5 - 21/4
= 56/20 - 105/20
= -49/20
-----------
Or
= 2.8 - 5.25
= -2.45